OpenStudy (anonymous):

Convert the decimal to a fraction. Show all work 0.55

OpenStudy (asnaseer):

\[0.55=\frac{55}{100}=\frac{11}{20}\]

OpenStudy (anonymous):

\[.55=\frac{55}{100}=\frac{5\times 11}{100}=\frac{11}{20}\]